Reports 1-1 of 1 Clear search Modify search
menzione, ciardelli, ballardin, berni - 11:47 Friday 24 September 2021 (53259) Print this report
TCS chiller BCK failure

This morning, the operator contacted TCS experts for an increasing of the temperature of the TCS chiller BCK that reached 30°C.
At 07:40 UTC, me and Marco went in TCS Chiller Room to check the status of the chiller BCK and we discovered that the cooling system was stucked.
We replace the BCK chiller (pict1) with the Spare one (pict2), re level the chiller with ultrapure water, restarted it, waited that everithing worked properly and the temperature stibilized at 19°C (pict3), restarted the process and the comunication with tne bridge.
Situation in standard state at 09:00 UTC.

The old chiller will be repaired asap. TCS experts informed.

Images attached to this report
Search Help